# S, T, U Parameters in The B-LSSM

### v1

@@ Abstract
Using the pinch technique, we compute the one-loop vertices of weak interactions in the B-LSSM and incorporate their pinch contributions into the gauge boson self-energies. Compared to the definitions of the $S$, $T$, and $U$ parameters in the Standard Model based on the $SU(2)_{L}\otimes U(1)_{Y}$ group, the corresponding parameters in the B-LSSM are modified. We provide these redefined $S$, $T$, and $U$ parameters and demonstrate the convergence of the results. In the framework of the low-energy effective Lagrangian for weak interactions, the $S$, $T$, and $U$ parameters can be expressed as functions of certain parameters in the B-LSSM. The updated experimental and fitting results constrain the parameter space of the B-LSSM strongly.
@@ I Introduction
The measurement results for the mass of the $W$ boson, as reported by the ATLAS experimental group in 2023 ATLAS:2023fsi , are in close agreement with the theoretical predictions of the Standard Model (SM). In comparison to the SM, the inclusion of the additional $U(1)_{B-L}$ gauge group in the B-LSSM leads to the emergence of a more substantial $Z^{\prime}$ boson, which in turn modifies the original definition of the Weinberg angle. The electroweak radiative corrections of observable quantities can be obtained through the oblique parameter method, and the superfields in the B-LSSM exert a notable influence on these oblique parameters. In this paper, we derive the gauge-invariant gauge boson self-energies in the B-LSSM. Basing on experimental results, we impose constraints on the parameters of the model.
@@
The $S$, $T$, and $U$ parameters Grimus:2008nb ; Maksymyk:1993zm ; Cacciapaglia:2004rb ; Lavoura:1992np ; Haywood:1999qg ; Cacciapaglia:2004jz ; Burgess:1993mg ; Asadi:2022xiy ; Long:1999bny ; Pich:2013fea present an extension of the method developed by Peskin Peskin:1991sw , building on the work of Kennedy and Lynn Kennedy:1988td , to handle radiative corrections in electroweak interaction processes based on the $SU(2)\times U(1)$ gauge group. This approach expresses radiative corrections to observables as a linear combination of the $S$, $T$, and $U$ parameters. Although the B-LSSM introduces an additional $U(1)_{B-L}$ gauge group compared to the SM, the definition of the $S$, $T$, and $U$ parameters remains valid at the electroweak scale. Because the effect of the new $U(1)_{B-L}$ gauge group on electroweak observables is negligible at tree level. The radiative corrections resulting from the new gauge group can be safely ignored.
@@
When calculating physical observables or considering physical processes, their gauge invariance is typically taken into account. It becomes critical when extracting physical information from S-matrix elements, as gauge invariance cannot always be guaranteed. The oblique parameters discussed in this paper must account for gauge invariance. To address this issue, the pinch technique (PT), a well-established method based on the background field method (BFM), is employed. This technique has been widely used in recent times. For more details, see references Binosi:2009qm ; Hashimoto:1994ct ; Denner:1994xt ; Denner:1994nn ; Denner:1996wn ; Denner:1995jd ; Papavassiliou:1993ex ; Degrassi:1992ue ; Binosi:2002ft ; Binosi:2003rr ; Papavassiliou:1995gs ; Binosi:2002ez ; Papavassiliou:1996zn ; Papavassiliou:1994pr .
@@
The following section outlines the structure of the paper. In Section II, we provide a brief introduction to the B-LSSM and the construction of the $Z^{{}^{\prime}}$ boson. In Section III, we present the gauge-invariant self-energies in the B-LSSM, derived using the pinch technique. In Section IV, we introduce the gauge-invariant $S$, $T$, and $U$ parameters in the B-LSSM, and based on these results, we further constrain the $s^{\prime}$ by incorporating effective Lagrangian techniques.In Section V, we proved the divergence cancellation of two sectors. In Section VI, we analyze the experimental data’s constraints on the parameter space. Finally, in Section VII, we present our conclusions.
@@ II The B-L SSM
$D_{\mu}=\partial_{\mu}-i\left(\begin{array}[]{cc}Y&Y_{B}\\ \end{array}\right)\left(\begin{array}[]{cc}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\\ \end{array}\right)\left(\begin{array}[]{c}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\\ \end{array}\right),$ (9)
where $\widetilde{A}_{\mu}^{Y}$ and $\widetilde{A}_{\mu}^{B}$ denote the gauge fields associated with the two $U(1)$ gauge groups. $Y$ and $Y_{B}$ denote the hypercharge and $B-L$ charge respectively. It is easier to work with non-canonical covariant derivatives instead of off-diagonal field-strength tensors:
$D_{\mu}=\partial_{\mu}-i\left(\begin{array}[]{cc}Y&Y_{B}\\ \end{array}\right)\left(\begin{array}[]{cc}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\\ \end{array}\right)R^{T}R\left(\begin{array}[]{c}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\\ \end{array}\right)$ (15)
@@
$\left(\begin{array}[]{cc}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\\ \end{array}\right)R^{T}=\left(\begin{array}[]{cc}g_{1}&g_{Y}\\ 0&g_{B}\\ \end{array}\right),\quad R\left(\begin{array}[]{c}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\\ \end{array}\right)=\left(\begin{array}[]{c}A_{\mu}^{Y}\\ A_{\mu}^{B}\\ \end{array}\right)$ (24)
@@
$M^{2}_{N}=\frac{1}{8}\left(\begin{array}[]{ccc}g_{1}^{2}v^{2}&g_{1}g_{2}v^{2}&g_{1}g_{Y}v^{2}\\ g_{1}g_{2}v^{2}&g_{2}^{2}v^{2}&g_{2}g_{Y}v^{2}\\ g_{1}g_{Y}v^{2}&g_{2}g_{Y}v^{2}&g_{Y}^{2}v^{2}+g_{B}^{2}u^{2}\\ \end{array}\right)$ (28)
@@
$R_{N}=\left(\begin{array}[]{ccc}c_{W}&-s_{W}&0\\ s_{W}&c_{W}&0\\ 0&0&1\\ \end{array}\right)\left(\begin{array}[]{ccc}1&0&0\\ 0&c^{\prime}&-s^{\prime}\\ 0&s^{\prime}&c^{\prime}\\ \end{array}\right)=\left(\begin{array}[]{ccc}c_{W}&-s_{W}c^{\prime}&s_{W}s^{\prime}\\ s_{W}&c_{W}c^{\prime}&-c_{W}s^{\prime}\\ 0&s^{\prime}&c^{\prime}\\ \end{array}\right)$ (38)
@@
$\left(\begin{array}[]{c}\gamma\\ Z\\ Z^{{}^{\prime}}\\ \end{array}\right)=R_{N}^{T}\left(\begin{array}[]{c}A^{Y}\\ B^{3}\\ A^{B}\\ \end{array}\right)$ (45)
@@
$\frac{1}{4}\left(\begin{array}[]{ccc}0&0&0\\ 0&G^{2}v^{2}&-Gg_{YB}v^{2}\\ 0&-Gg_{YB}v^{2}&g_{YB}^{2}v^{2}+4g_{B}^{2}u^{2}\\ \end{array}\right)$
$=\left(\begin{array}[]{ccc}1&0&0\\ 0&c^{\prime}&-s^{\prime}\\ 0&s^{\prime}&c^{\prime}\\ \end{array}\right)\left(\begin{array}[]{ccc}0&0&0\\ 0&m_{Z}^{2}&0\\ 0&0&m_{Z^{\prime}}^{2}\\ \end{array}\right)\left(\begin{array}[]{ccc}1&0&0\\ 0&c^{\prime}&s^{\prime}\\ 0&-s^{\prime}&c^{\prime}\\ \end{array}\right)$ (57)
$=\left(\begin{array}[]{ccc}0&0&0\\ 0&{c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}&c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})\\ 0&c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})&{s^{\prime}}^{2}m^{2}_{Z}+{c^{\prime}}^{2}m^{2}_{Z^{\prime}}\\ \end{array}\right)$ (61)
Where $G$ is Electric weak coupling coefficient $G=\frac{e}{s_{W}c_{W}}$. Then, $g_{B}$ and $g_{Y}B$ can be represented by some observable measurements:
$g_{YB}$ $=G\frac{c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}}$
$g_{B}$ $=\frac{G}{2}\frac{xm_{Z}m_{Z^{\prime}}}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}}$ (62)
@@ III Pinch Technique
The gauge-invariant self energys of $Z$ and $Z^{\prime}$ can be obtained in any neutral flow process. Following the approach of Papavasiliou Papavassiliou:1994pr , consider the scattering process $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$, The $S$-matrix element of this process should remain independent of the gauge-fixing parameters $\xi_{i}$ at any order of perturbative calculation. The propagators for the gauge bosons $\gamma$, $Z$ and $Z^{{}^{\prime}}$ which include the gauge-fixing parameters, are given by:
$\Delta_{\mu\nu}{\gamma}(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{\gamma})\frac{q_{\mu}q_{\nu}}{q^{2}}]$ (63)
$\Delta_{\mu\nu}Z(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{Z})\frac{q_{\mu}q_{\nu}}{q^{2}-\xi_{Z}m^{2}_{Z}}]$ (64)
$\Delta_{\mu\nu}Z^{{}^{\prime}}(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{Z^{{}^{\prime}}})\frac{q_{\mu}q_{\nu}}{q^{2}-\xi_{Z^{{}^{\prime}}}m^{2}_{Z^{{}^{\prime}}}}]$ (65)
@@
Figure 1: Graphs (a)-(c) are some of the 1-loop contributions to the S-matrx of this four-fermion process. The pinch parts of (b) and (c), shown in (e) and (f) are absorbed into (d), making it become a gauge-invariant effective self-energy.
@@
$=S_{i}^{-1}(p+k)P_{L}-S_{j}^{-1}(p)P_{R}+m_{i}P_{L}-m_{j}P_{R}$
Here, $P_{R,L}=\frac{1\pm\gamma_{5}}{2}$. The terms $S_{i}^{-1}(p+k)$ and $S_{j}^{-1}(p)$ vanish on shell, leaving behind $m_{i}P_{L}-m_{j}P_{R}$. When charged $W^{\pm}$ couples to fermions, the masses $m_{i}$ and $m_{j}$ differ. For neutral $Z$ fermions of equal mass,$m_{i}=m_{j}$.
@@
$\Gamma^{P\xi}_{\mu\nu\alpha}=\frac{1}{\xi}[(q+k)_{\nu}g_{\mu\alpha}+k_{\mu}g_{\nu\alpha}]$ (71)
$\Gamma^{\xi}_{\mu\nu\alpha}=2q_{\mu}g_{\nu\alpha}-2q_{\nu}g_{\mu\alpha}-(2k+q)_{\alpha}g_{\mu\nu}+(1-\frac{1}{\xi})[k_{\nu}g_{\mu\alpha}+(k+q)_{\mu}g_{\nu\alpha}],$
@@
$q^{\alpha}\Gamma^{\xi}_{\mu\nu\alpha}=\Delta^{-1}_{\mu\nu}(k,\xi)-\Delta^{-1}_{\mu\nu}(k+q,\xi)$ (72)
@@
$\Delta^{i}_{\mu\nu}(q)=\mathscr{D}^{i}_{\mu\nu}(q)-\frac{q_{\mu}q_{\nu}}{M^{2}_{i}}D^{i}(q)$ (73)
@@
$D^{i}(q,\xi)=\frac{i}{q^{2}-\xi_{i}M_{i}^{2}}$
$\mathscr{D}^{i}_{\mu\nu}(q)=[g_{\mu\nu}-\frac{q_{\mu}q_{\nu}}{M^{2}_{i}}]\frac{-i}{q^{2}-M_{i}^{2}}$ (74)
@@
$=$ $\Delta^{i}_{\nu\mu}(q,\xi_{i})\mathscr{D}_{i}^{-1}(q)^{\mu\alpha}-iq_{\nu}q^{\alpha}D_{i}(q,\xi)$
$q_{\mu}=$ $-i\{q^{2}D_{i}(q,\xi_{i})q_{\mu}+M_{i}^{2}q^{\nu}\Delta^{i}_{\nu\mu}(q,\xi_{i})\}$ (75)
@@
$\mathscr{D}_{i}^{-1}(q)^{\mu\alpha}=i\{g^{\mu\alpha}(q^{2}-M_{i}^{2})-q^{\mu}q^{\alpha}\}$ (76)
@@
${\Delta^{\gamma}_{\mu\nu}}^{-1}(q)$ $=iq^{2}[g^{\mu\nu}+(1-\frac{1}{\xi_{\gamma}})\frac{q_{\mu}q_{\nu}}{q^{2}}]$ (77)
@@
$\Pi^{P}_{WW}(q^{2})=-4g_{2}^{2}(q^{2}-M_{W}^{2})\{s_{W}^{2}I_{W\gamma}(q^{2})+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(q^{2})+{s^{\prime}}^{2}I_{WZ}(q^{2}))\}$ (78)
@@
$\frac{g_{2}}{2}\gamma^{\mu}P_{L}=(-i)s_{W}(e\gamma e)^{\mu}-ic_{W}c^{{}^{\prime}}(eZe)^{\mu}+ic_{W}s^{{}^{\prime}}(eZ^{{}^{\prime}}e)^{\mu}$ (80)
By combining Eq. (75) and Eq. (80), the following results can be obtained:
$\Pi^{P}_{\gamma\gamma}(q^{2})=-4g_{2}^{2}s_{W}^{2}q^{2}I_{WW}(q^{2})$
$\Pi^{P}_{\gamma Z}(q^{2})=-2g_{2}^{2}s_{W}c_{W}c^{{}^{\prime}}(2q^{2}-M_{Z}^{2})I_{WW}(q^{2})$
$\Pi^{P}_{\gamma Z^{{}^{\prime}}}(q^{2})=-2g_{2}^{2}s_{W}c_{W}s^{{}^{\prime}}(2q^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2})$
$\Pi^{P}_{ZZ}(q^{2})=-4g_{2}^{2}c^{2}_{W}{c^{{}^{\prime}}}^{2}(q^{2}-M_{Z}^{2})I_{WW}(q^{2})$
$\Pi^{P}_{Z^{{}^{\prime}}Z^{{}^{\prime}}}(q^{2})=-4g_{2}^{2}c^{2}_{W}{s^{{}^{\prime}}}^{2}(q^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2})$
$\Pi^{P}_{ZZ^{{}^{\prime}}}(q^{2})=2g_{2}^{2}c^{2}_{W}s^{{}^{\prime}}c^{{}^{\prime}}(2q^{2}-M_{Z}^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2})$ (81)
[…]

### v2

@@ Abstract
Using the pinch technique, we compute the one-loop vertices of weak interactions in the B-LSSM and incorporate their pinch contributions into the gauge boson self-energies. Compared to the definitions of the $S$, $T$, and $U$ parameters in the Standard Model based on the $SU(2)_{L}\otimes U(1)_{Y}$ group, the corresponding parameters in the local B-L gauge symmetry (B-LSSM) are modified. We provide these redefined $S$, $T$, and $U$ parameters and demonstrate the convergence of the results. In the framework of the low-energy effective Lagrangian for weak interactions, the $S$, $T$, and $U$ parameters can be expressed as functions of certain parameters in the B-LSSM. The updated experimental and fitting results constrain the parameter space of the B-LSSM strongly.
@@ I Introduction
The measurement results for the mass of the $W$ boson, as reported by the ATLAS experimental group in 2023 ATLAS:2023fsi , are in close agreement with the theoretical predictions of the Standard Model (SM). This high-precision agreement provides a stringent benchmark for probing theories beyond the SM. In particular, models with extended gauge sectors, such as the B-LSSM, must accommodate this result, which can lead to significant constraints on their parameter space. In comparison to the SM, the inclusion of the additional $U(1)_{B-L}$ gauge group in the B-LSSM leads to the emergence of a more substantial $Z^{\prime}$ boson, which in turn modifies the original definition of the Weinberg angle. The electroweak radiative corrections of observable quantities can be obtained through the oblique parameter method, and the superfields in the B-LSSM exert a notable influence on these oblique parameters. In this paper, we derive the gauge-invariant gauge boson self-energies in the B-LSSM. Based on experimental results, we impose constraints on the parameters of the model.
@@
The $S$, $T$, and $U$ parameters Grimus:2008nb ; Maksymyk:1993zm ; Cacciapaglia:2004rb ; Lavoura:1992np ; Haywood:1999qg ; Cacciapaglia:2004jz ; Burgess:1993mg ; Asadi:2022xiy ; Long:1999bny ; Pich:2013fea ; Rehman:2025urc present an extension of the method developed by Peskin Peskin:1991sw , building on the work of Kennedy and Lynn Kennedy:1988td , to handle radiative corrections in electroweak interaction processes. Although the B-LSSM introduces an additional $U(1)_{B-L}$ gauge group compared to the SM, the oblique parameter framework remains a powerful tool at the electroweak scale. The new gauge sector modifies the observables directly via tree-level gauge kinetic mixing and mass matrix deformations, which can be elegantly encoded into the effective $S$, $T$, and $U$ parameters.
@@
When calculating physical observables or considering physical processes, their gauge invariance is typically taken into account. It becomes critical when extracting physical information from S-matrix elements, as gauge invariance cannot always be guaranteed. The oblique parameters discussed in this paper must account for gauge invariance. To address this issue, the pinch technique (PT), a well-established method based on the background field method (BFM), is employed. This technique has found widespread adoption and application. For more details, see references Binosi:2009qm ; Hashimoto:1994ct ; Denner:1994xt ; Denner:1994nn ; Denner:1996wn ; Denner:1995jd ; Papavassiliou:1993ex ; Degrassi:1992ue ; Binosi:2002ft ; Binosi:2003rr ; Papavassiliou:1995gs ; Binosi:2002ez ; Papavassiliou:1996zn ; Papavassiliou:1994pr .
While many studies on oblique parameters rely heavily on Standard Model Effective Field Theory (SMEFT) global fits, it is important to recognize the conceptual distinction between these effective coefficients and the full theoretical calculations. To provide a more fundamental theoretical foundation, this work aims to avoid ”over-fitting” the full theory directly to these effective parameters. Instead, we start from the formal theory of the B-LSSM to rigorously derive the gauge-invariant analytic expressions at the one-loop level, focusing on the physical trends and sensitivities revealed by the theoretical calculation itself.
Based on this robust framework, our numerical analysis strategy focuses on characterizing the dominant new physics sources. First, we systematically analyze how the mass mixing in the key supersymmetric sectors (namely the squark and neutralino-chargino sectors) influences the electroweak precision observables at the loop level. Second, for the additional $U(1)_{B-L}$ gauge parameters, we employ a complementary approach, utilizing the effective Lagrangian to impose direct constraints at the tree level. This dual approach lays the necessary methodological foundation and serves as a physical template for understanding how different symmetry-breaking sectors collectively impact low-energy precision measurements.
@@
The structure of this paper is outlined as follows: In Section II, we provide a brief introduction to the B-LSSM and the construction of the $Z^{{}^{\prime}}$ boson. In Section III, we present the gauge-invariant self-energies in the B-LSSM, derived using the pinch technique. In Section IV, we introduce the gauge-invariant $S$, $T$, and $U$ parameters in the B-LSSM, and based on these results, we further constrain the gauge mixing angle $s^{\prime}$ by incorporating effective Lagrangian techniques.In Section V, we proved the divergence cancellation of two sectors. In Section VI, we analyze the experimental data’s constraints on the parameter space. Finally, in Section VII, we present our conclusions.
@@ II The B-L SSM
$D_{\mu}=\partial_{\mu}-i\begin{pmatrix}Y&Y_{B}\end{pmatrix}\begin{pmatrix}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\end{pmatrix}\begin{pmatrix}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\end{pmatrix}$, (4)
where $\widetilde{A}_{\mu}^{Y}$ and $\widetilde{A}_{\mu}^{B}$ denote the gauge fields associated with the two $U(1)$ gauge groups. $Y$ and $Y_{B}$ denote the hypercharge and BLB-L charge respectively. It is easier to work with non-canonical covariant derivatives instead of off-diagonal field-strength tensors:
$D_{\mu}=\partial_{\mu}-i\begin{pmatrix}Y&Y_{B}\end{pmatrix}\begin{pmatrix}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\end{pmatrix}R^{T}R\begin{pmatrix}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\end{pmatrix}$. (5)
@@
$\begin{pmatrix}\widetilde{g}&\widetilde{g}_{YB}\\ \widetilde{g}_{BY}&\widetilde{g}_{B}\end{pmatrix}R^{T}=\begin{pmatrix}g_{1}&g_{Y}\\ 0&g_{B}\end{pmatrix},\quad R\begin{pmatrix}\widetilde{A}_{\mu}^{Y}\\ \widetilde{A}_{\mu}^{B}\end{pmatrix}=\begin{pmatrix}A_{\mu}^{Y}\\ A_{\mu}^{B}\end{pmatrix}$. (6)
@@
$M^{2}_{N}=\frac{1}{8}\begin{pmatrix}g_{1}^{2}v^{2}&g_{1}g_{2}v^{2}&g_{1}g_{Y}v^{2}\\ g_{1}g_{2}v^{2}&g_{2}^{2}v^{2}&g_{2}g_{Y}v^{2}\\ g_{1}g_{Y}v^{2}&g_{2}g_{Y}v^{2}&g_{Y}^{2}v^{2}+g_{B}^{2}u^{2}\end{pmatrix}$. (7)
@@
$R_{N}=\begin{pmatrix}c_{W}&-s_{W}&0\\ s_{W}&c_{W}&0\\ 0&0&1\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c^{\prime}&-s^{\prime}\\ 0&s^{\prime}&c^{\prime}\end{pmatrix}=\begin{pmatrix}c_{W}&-s_{W}c^{\prime}&s_{W}s^{\prime}\\ s_{W}&c_{W}c^{\prime}&-c_{W}s^{\prime}\\ 0&s^{\prime}&c^{\prime}\end{pmatrix}$, (8)
@@
$\begin{pmatrix}\gamma\\ Z\\ Z^{\prime}\end{pmatrix}=R_{N}^{T}\begin{pmatrix}A^{Y}\\ B^{3}\\ A^{B}\end{pmatrix}$. (9)
@@
$\frac{1}{4}\begin{pmatrix}0&0&0\\ 0&G^{2}v^{2}&-Gg_{YB}v^{2}\\ 0&-Gg_{YB}v^{2}&g_{YB}^{2}v^{2}+4g_{B}^{2}u^{2}\end{pmatrix}$
$=\begin{pmatrix}1&0&0\\ 0&c^{\prime}&-s^{\prime}\\ 0&s^{\prime}&c^{\prime}\end{pmatrix}\begin{pmatrix}0&0&0\\ 0&m_{Z}^{2}&0\\ 0&0&m_{Z^{\prime}}^{2}\end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c^{\prime}&s^{\prime}\\ 0&-s^{\prime}&c^{\prime}\end{pmatrix}$
$=\begin{pmatrix}0&0&0\\ 0&{c^{\prime}}^{2}m_{Z}^{2}+{s^{\prime}}^{2}m_{Z^{\prime}}^{2}&c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})\\ 0&c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})&{s^{\prime}}^{2}m^{2}_{Z}+{c^{\prime}}^{2}m^{2}_{Z^{\prime}}\end{pmatrix}$, (10)
where $G$ is electroweak coupling coefficient $G=\frac{e}{s_{W}c_{W}}$, $v^{2}=v_{1}^{2}+v_{2}^{2}$ and $u^{2}=u_{1}^{2}+u_{2}^{2}$. Then, $g_{B}$ and $g_{YB}$ can be represented by some observable measurements:
$g_{YB}$ $=G\frac{c^{\prime}s^{\prime}(m^{2}_{Z}-m^{2}_{Z^{\prime}})}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}},$
$g_{B}$ $=\frac{G}{2}\frac{xm_{Z}m_{Z^{\prime}}}{{c^{\prime}}^{2}m^{2}_{Z}+{s^{\prime}}^{2}m^{2}_{Z^{\prime}}}.$ (11)
@@ III Pinch Technique
The gauge-invariant self-energies of the $Z$ and $Z^{\prime}$ bosons can be extracted from any neutral-current process. Following the approach of Papavasiliou Papavassiliou:1994pr , we consider the scattering process $e^{+}e^{-}\rightarrow\mu^{+}\mu^{-}$, The $S$-matrix element of this process should remain independent of the gauge-fixing parameters $\xi_{i}$ at any order of perturbative calculation. The propagators for the gauge bosons $\gamma$, $Z$ and $Z^{{}^{\prime}}$ which include the gauge-fixing parameters, are given by:
$\Delta_{\mu\nu}{\gamma}(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{\gamma})\frac{q_{\mu}q_{\nu}}{q^{2}}],$ (12)
$\Delta_{\mu\nu}Z(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{Z})\frac{q_{\mu}q_{\nu}}{q^{2}-\xi_{Z}m^{2}_{Z}}],$ (13)
$\Delta_{\mu\nu}Z^{{}^{\prime}}(q)=\frac{-i}{q^{2}}[g_{\mu\nu}-(1-\xi_{Z^{{}^{\prime}}})\frac{q_{\mu}q_{\nu}}{q^{2}-\xi_{Z^{{}^{\prime}}}m^{2}_{Z^{{}^{\prime}}}}].$ (14)
@@
Figure 1: Diagram (a)-(c) show representative one-loop contributions to the S-matrix of the four-fermion process. Diagram (a) is the conventional gauge-dependent self-energy. The pinch parts extracted from diagrams (b) and (c)–shown separately in (e) and (f)–are systematically absorbed into (a), thereby converting it into the gauge-invariant effective self-energy displayed in (d).
@@
$=S_{i}^{-1}(p+k)P_{L}-S_{j}^{-1}(p)P_{R}+m_{i}P_{L}-m_{j}P_{R}.$
Here, $P_{R,L}=\frac{1\pm\gamma_{5}}{2}$. The terms $S_{i}^{-1}(p+k)$ and $S_{j}^{-1}(p)$ vanish on shell, leaving behind $m_{i}P_{L}-m_{j}P_{R}$. When charged $W^{\pm}$ couples to fermions, the masses $m_{i}$ and $m_{j}$ differ. For the $Z$ boson coupling to a fermion-antifermion pair of equal mass, $m_{i}=m_{j}$.
@@
$\Gamma^{P\xi}_{\mu\nu\alpha}=\frac{1}{\xi}[(q+k)_{\nu}g_{\mu\alpha}+k_{\mu}g_{\nu\alpha}],$ (20)
$\Gamma^{\xi}_{\mu\nu\alpha}=2q_{\mu}g_{\nu\alpha}-2q_{\nu}g_{\mu\alpha}-(2k+q)_{\alpha}g_{\mu\nu}+(1-\frac{1}{\xi})[k_{\nu}g_{\mu\alpha}+(k+q)_{\mu}g_{\nu\alpha}].$
@@
$q^{\alpha}\Gamma^{\xi}_{\mu\nu\alpha}=\Delta^{-1}_{\mu\nu}(k,\xi)-\Delta^{-1}_{\mu\nu}(k+q,\xi).$ (21)
@@
$\Delta^{i}_{\mu\nu}(q)=\mathscr{D}^{i}_{\mu\nu}(q)-\frac{q_{\mu}q_{\nu}}{M^{2}_{i}}D^{i}(q),$ (22)
@@
$D^{i}(q,\xi)=\frac{i}{q^{2}-\xi_{i}M_{i}^{2}},$
$\mathscr{D}^{i}_{\mu\nu}(q)=[g_{\mu\nu}-\frac{q_{\mu}q_{\nu}}{M^{2}_{i}}]\frac{-i}{q^{2}-M_{i}^{2}}.$ (23)
@@
$=$ $\Delta^{i}_{\nu\mu}(q,\xi_{i})\mathscr{D}_{i}^{-1}(q)^{\mu\alpha}-iq_{\nu}q^{\alpha}D_{i}(q,\xi),$
$q_{\mu}=$ $-i\{q^{2}D_{i}(q,\xi_{i})q_{\mu}+M_{i}^{2}q^{\nu}\Delta^{i}_{\nu\mu}(q,\xi_{i})\},$ (24)
@@
$\mathscr{D}_{i}^{-1}(q)^{\mu\alpha}=i\{g^{\mu\alpha}(q^{2}-M_{i}^{2})-q^{\mu}q^{\alpha}\}.$ (25)
@@
${\Delta^{\gamma}_{\mu\nu}}^{-1}(q)$ $=iq^{2}[g^{\mu\nu}+(1-\frac{1}{\xi_{\gamma}})\frac{q_{\mu}q_{\nu}}{q^{2}}].$ (26)
@@
$\Pi^{P}_{WW}(q^{2})=-4g_{2}^{2}(q^{2}-M_{W}^{2})\{s_{W}^{2}I_{W\gamma}(q^{2})+c_{W}^{2}({c^{\prime}}^{2}I_{WZ}(q^{2})+{s^{\prime}}^{2}I_{WZ^{\prime}}(q^{2}))\}.$ (27)
@@
$\frac{g_{2}}{2}\gamma^{\mu}P_{L}=(-i)s_{W}(e\gamma e)^{\mu}-ic_{W}c^{\prime}(eZe)^{\mu}+ic_{W}s^{\prime}(eZ^{\prime}e)^{\mu},$ (29)
Where $(eZe)$ and $(eZ^{\prime}e)$ denote the vertex of the corresponding particles. By combining Eq. (24) and Eq. (29), the following results can be obtained:
$\Pi^{P}_{\gamma\gamma}(q^{2})=-4g_{2}^{2}s_{W}^{2}q^{2}I_{WW}(q^{2}),$
$\Pi^{P}_{\gamma Z}(q^{2})=-2g_{2}^{2}s_{W}c_{W}c^{{}^{\prime}}(2q^{2}-M_{Z}^{2})I_{WW}(q^{2}),$
$\Pi^{P}_{\gamma Z^{{}^{\prime}}}(q^{2})=-2g_{2}^{2}s_{W}c_{W}s^{{}^{\prime}}(2q^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2}),$
$\Pi^{P}_{ZZ}(q^{2})=-4g_{2}^{2}c^{2}_{W}{c^{{}^{\prime}}}^{2}(q^{2}-M_{Z}^{2})I_{WW}(q^{2}),$
$\Pi^{P}_{Z^{{}^{\prime}}Z^{{}^{\prime}}}(q^{2})=-4g_{2}^{2}c^{2}_{W}{s^{{}^{\prime}}}^{2}(q^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2}),$
$\Pi^{P}_{ZZ^{{}^{\prime}}}(q^{2})=2g_{2}^{2}c^{2}_{W}s^{{}^{\prime}}c^{{}^{\prime}}(2q^{2}-M_{Z}^{2}-M_{Z^{{}^{\prime}}}^{2})I_{WW}(q^{2}).$ (30)
By utilizing PT, we successfully derive gauge-invariant self-energies for the $W^{\pm}$, $Z$, $Z^{\prime}$ and $\gamma$ bosons, which are essential for subsequent calculations.